3x+y =-6
4x +z=0
2y+z=-2

Answer by Alan3354(69443)

3x+y =-6
4x +z=0
2y+z=-2
-------------
Substitution will work.
z = -4x
2y+z = -2
-4x+ 2y = -2
3x + y = -6 x2 --> 6x + 2y = -12
6x + 2y = -12
-4x+ 2y = -2
--------------- Subtract
10x = -10
x = -1
------
z = 4
----
y = -3